OpenStudy (anonymous):

A circle has a radius of 6 in. The circumscribed equilateral triangle will have an area of:

OpenStudy (anonymous):

was the answer wrong?

OpenStudy (anonymous):

No this is a different problem, ur answer was right ty for that

OpenStudy (anonymous):

U may use trig values of stndard angkes to know the length of each side

OpenStudy (anonymous):

I don't know any trig

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The height of the equilateral triangle will be 32×6=9 Length of the side =9×23√=63√ Area =1263√×9=273√in2

OpenStudy (anonymous):

Alright ty

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

http://openstudy.com/study#/updates/4fc819dfe4b022f1e12f984f this should help u

OpenStudy (anonymous):

Thanks everyone, u guys r a big help, I have to go now see ya guys later

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

The height would then be the longer leg of a 30-60-90 triangle, 6sqrt(3)*sqrt(3) = 6*3 = 18 A = 1/2 b h = 1/2 18 * 12sqrt(3) = 9 * 12sqrt(3) = 108sqrt(3)

ganeshie8 (ganeshie8):

that may avoid using trig directly, height = 18, use 30-60-90 special triangle to figure out the base

OpenStudy (anonymous):

thank you for the medal

OpenStudy (anonymous):

Area =1083√=12×3×3×3√